The manufacturing costs of Rosenthal Industries for the first three months of the year follow: Total Costs Production January $106,020 2,170 units February 132,760 3,270 March 164,920 5,270 Using the high-low method, determine (a) the variable cost per unit and (b) the total fixed cost. a. Variable cost per unit $ b. Total fixed cost $
Answer #1

a. Variable cost per unit = (164920-106020)/(5270-2170) = $19 per unit

b. Fixed cost = 106020 - 2170*19 = $6470